[Hallc_running] [New Logentry] RC Report -- 09/07/22

dhamilto at jlab.org dhamilto at jlab.org
Wed Sep 7 09:00:02 EDT 2022

Logentry Text:
Beam is currently off for the maintenance day. RadCon are coming over around 9am to perform a sweep, although there is not much planned in the hall today. If possible, Ops plan to restore beam before Swing shift. The SHMS is already at 10 degrees for the next kinematic setting.

We had a good day of data taking for the Q2=3.85, W=2.62, Th_SHMS=8.00 setting at 12 uA on LD2, with the only noteworthy issues being several HMS Q1 interlocks yesterday (see eg https://logbooks.jlab.org/entry/4038521). It came right back up after reset each time.

The plan for the rest of the week includes:
1) Low current BCM calibration using the Faraday cup when beam is restored today (coordinate with Dave Mack).
2) Take data at 38 uA on LD2 for the Q2=3.85, W=2.62, Th_SHMS=10.00 setting for around two shifts (including 60 mC of dummy target data).
3) Change back to the Q2=3.85, W=2.62, Th_SHMS=8.00 setting to take data at 12 uA on LD2 until Friday morning (including 200 mC of dummy target data).

Let's have a short meeting today at 11am to discuss.
Meeting ID: 160 710 0885 - Passcode: 595071


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/4038707
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20220907/4d9480ec/attachment.html>

More information about the Hallc_running mailing list